A car is travelling at a constant speed of 120 m/s. Passes a trooper hidden behind a Billboard one second of the car passes the Billboard, the trooper sets out from the Billboard to catch it, acelerating at a constant speed of 5 m/s².how long does it take her to take the car?
1
Expert's answer
2020-02-17T09:14:09-0500
"d=v(t+1)"

"d=0.5at^2"

Thus,


"v(t+1)=0.5at^2"

"120(t+1)=0.5(5)t^2"

"t=49\ s"